OpenStudy (anonymous):

Evaluate the integral ((e^t+2)/(e^t+2t))dt

OpenStudy (anonymous):

\[-\frac{1}{2}e^{2(1-t)}+K\]

OpenStudy (anonymous):

does it matter is t is greater than 0

OpenStudy (anonymous):

just to make sure I understand the problem: you asked \[\int\frac{e^{t+2}}{e^{t+2t}}\]?

OpenStudy (anonymous):

or \[\int\frac{e^t+2}{e^t+2t}dt\]

OpenStudy (anonymous):

because I answered the first , the second would be \[\ln\mid{e^t+2t}\mid+K\]

OpenStudy (anonymous):

the second it does matter that \[t>0\]

OpenStudy (anonymous):

ok